OpenStudy (anonymous):

Inflection points?: f(x)= (3/4)x^4 - x^3 - 3x^2 + 6x

OpenStudy (anonymous):

1. Find the inflection points. Show why there are exactly two inflection points for this function.

hartnn (hartnn):

to get inflection points, we equate 2nd derivative of the function to 0 f''(x) =0 so can you first find f''(x) ??

OpenStudy (anonymous):

F’’(x) = 9x^2 – 6x + 6

hartnn (hartnn):

+6 ? or -6 ?`

OpenStudy (anonymous):

Oh, sorry! f''(x) = 9x^2-6x-6

hartnn (hartnn):

yes, now solve 9x^2-6x-6=0

OpenStudy (anonymous):

I know that we have to factor 9x^2-6x-6, but I'm not very sure how to factor it. Another person had helped me with a similar question the other day and he had group factored it. Not sure which way to go?

OpenStudy (anonymous):

When I factor it, I get 3(3x^2 - 2x -2)

hartnn (hartnn):

you'll need quadratic formula know what that is ? (solutions are irrational)

OpenStudy (anonymous):

Yes, I do. okay, give me one quick moment to plug in my numbers.

hartnn (hartnn):

sure, take your time :)

OpenStudy (anonymous):

6 +/- (sqrt(-6^2 – 4(9*-6)))/2(9) =6 (sqrt(36 – 216) / 18) = 4.47

OpenStudy (anonymous):

is this correct?

hartnn (hartnn):

naah... b^2- 4ac = (-6)^2 - 4 (9)(-6) = 36 + 216 ....

OpenStudy (anonymous):

Yea, realized I was way off. I'm going to try and solve it in a different way. one sec.

hartnn (hartnn):

x = 1/3 – 1/3sqrt(7) x = 1/3 + 1/3sqrt(7)

hartnn (hartnn):

YES

OpenStudy (anonymous):

Okay, lol. thank you.

hartnn (hartnn):

welcome ^_^